48. The formula v=2sqrt6L is used by police to estimate the speed of car, V, in miles per hour based on the length of its skid marks (L, in feet) on dry pavement. How far will a car skid at a speed of 72 miles per hour?
Answer by checkley75(3666)

72=2sqrt6*L
72/2=sqrt6L
36=sqrt6L
36^2=6L
1296=6L
L=1296/6
L=216 estimated length of the skid @ 72 MPH.